# The Cost of Dynamic Reasoning: Demystifying AI Agents and Test-Time Scaling from an AI Infrastructure Perspective

### v1

@@ VIII Conclusion
This paper provides the first system-level characterization of AI agents from an AI infrastructure perspective. While these LLM-based agents demonstrate powerful reasoning capabilities through dynamic, multi-step interactions, they also introduce substantial compute, memory, and energy overheads that are orders of magnitude higher than conventional single-turn LLM inference. Our analysis shows that common agent design patterns, such as iterative reflection or parallel reasoning, incur heavy latency penalties and infrastructure costs, especially when deployed at scale. Moreover, test-time scaling yields sharply diminishing returns in accuracy, challenging the cost-effectiveness of current agent implementations.

### v2

@@ IX Conclusion
This paper provides the first system-level characterization of AI agents from an AI infrastructure perspective. While these LLM-based agents demonstrate powerful reasoning capabilities, they also introduce substantial energy overheads that are orders of magnitude higher than conventional single-turn LLM inference. Our analysis shows that common agent design patterns incur heavy latency penalties and infrastructure costs, especially when deployed at scale. Moreover, test-time scaling yields sharply diminishing returns in accuracy, challenging the cost-effectiveness of current agent implementations.
@@
Acknowledgment
This work was partly supported by Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.RS-2024-00438851, (SW Starlab) High-performance Privacy-preserving Machine Learning System and System Software), (No.RS-2024-00395134, DPU-Centric Datacenter Architecture for Next-Generation AI Devices), (No.RS-2025-02264029, Implementation and Validation of an AI Semiconductor-Based Data Center Composable Cluster Infrastructure, 30%), and by Samsung Research Funding Center of Samsung Electronics (SRFC-IT2402-03). Minsoo Rhu is the corresponding author.